OpenStudy (anonymous):

So confused on this question.. I keep getting getting 65 = 13 how can 65 be equal to 13 someone care to explain?

OpenStudy (anonymous):

oops forgot to include the question n/5 = 13

OpenStudy (anonymous):

Explain how u do it lol and i will reply

OpenStudy (anonymous):

n/5 = 13 n*5 =13 65 = 13 o.O

OpenStudy (anonymous):

n/5 = 13 n=13*5 Why n*5?? and for that matter n*5=5n n=65 and anyways u don't have to do it here.! Welcome to openstudy!

OpenStudy (anonymous):

Thanks you're awesome I'm stuck on another one if you don't mind i keep getting stuck at the division.. it's this y/4 - 3 = -4

OpenStudy (anonymous):

Yes my pleasure

OpenStudy (anonymous):

Give me a second!!!!

OpenStudy (anonymous):

\[\frac{ y }{ 4 }-3=-4 \]

OpenStudy (anonymous):

Well that was the question lol

OpenStudy (anonymous):

It's easy don't panic lmao

OpenStudy (anonymous):

Hey rather than me explaining it take a look at this for the answer scroll it down

OpenStudy (anonymous):

And thanks for the medal bro!!

OpenStudy (anonymous):

Thanks for the help :D